OpenStudy (anonymous):

A regular pentagon is a shape with 5 sides of equal length. Look at the relationship between the length of a side of a pentagon and the perimeter of the pentagon. Pentagon Side Pentagon Perimeter 2 ft 10 ft 4 ft 20 ft 6 ft 30 ft 8 ft 40 ft Which equation can be used to find p, the perimeter of the pentagon with sides of length s?

OpenStudy (anonymous):

p = (1/5) + s...p = s + s + s + s + s + s...p = 5 × s... or p = (1/5) × s

OpenStudy (anonymous):